A 20 g ball of clay traveling east at 3.6 m/s collides with a 30 g ball of clay traveling north at 2.0 m/s. What are the speed and the direction of the resulting 50 g ball of clay?
Answer #1
Concepts and reason

The concept required to solve this problem is Conservation of momentum.

Initially, calculate the momentum of each clay ball before the collision and add them to get the total initial momentum in horizontal and vertical direction.

Later, use the conservation of momentum in both the directions to solve for final velocity of the clay. Then use the magnitude equation to solve for the speed of the clay.

Finally, use the tangent function to solve for the direction of final velocity of the clay.

Fundamentals

The momentum of an object is the product of its mass and velocity. The magnitude of momentum is expressed as,

p=mvp = mv

Here, pp is the momentum, mm is the mass and vv is the speed.

The conservation of momentum principle says that total momentum of an isolated system is always conserved. This can be expressed as,

pinitial=pfinal{p_{{\rm{initial}}}} = {p_{{\rm{final}}}}

Here, pinitial{p_{{\rm{initial}}}} is the total initial momentum of the system and pfinal{p_{{\rm{final}}}} is the total final momentum.

The speed is the magnitude of the velocity that is,

v=vx2+vy2v = \sqrt {v_x^2 + v_y^2}

Here, vx{v_x} is the velocity in horizontal direction and vy{v_y} is the velocity in the vertical direction.

The direction of the velocity is given by inverse of tangent function as,

θ=arctan(vyvx)\theta = \arctan \left( {\frac{{{v_y}}}{{{v_x}}}} \right)

Here, vx{v_x} is the velocity in horizontal direction and vy{v_y} is the velocity in the vertical direction.

The expression for momentum is,

P=mvP = mv

Substitute p1ix{p_{1{\rm{ix}}}} for pp, 20g20{\rm{ g}} for mm, and 3.6m/s3.6{\rm{ m/s}} for vv in the above equation to calculate the initial momentum of the clay 1, p1ix{p_{1{\rm{ix}}}} in horizontal direction.

p1ix=(20g)(3.6m/s)=(20g(103kg1g))(3.6m/s)=0.072kgm/s\begin{array}{c}\\{p_{1{\rm{ix}}}} = \left( {{\rm{20 g}}} \right)\left( {3.6{\rm{ m/s}}} \right)\\\\ = \left( {{\rm{20 g}}\left( {\frac{{{{10}^{ - 3}}{\rm{ kg}}}}{{1{\rm{ g}}}}} \right)} \right)\left( {3.6{\rm{ m/s}}} \right)\\\\ = 0.072{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

Substitute p1iy{p_{1{\rm{iy}}}} for pp, 20g20{\rm{ g}} for mm, and 0m/s0{\rm{ m/s}} for vv in the equation p=mvp = mv to calculate the initial momentum of the clay 1, p1iy{p_{1{\rm{iy}}}} in vertical direction.

p1iy=(20g)(0m/s)=0kgm/s\begin{array}{c}\\{p_{1{\rm{iy}}}} = \left( {{\rm{20 g}}} \right)\left( {0{\rm{ m/s}}} \right)\\\\ = 0{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

Substitute p2ix{p_{{\rm{2ix}}}} for pp, 30g{\rm{30 g}} for mm, and 0m/s0{\rm{ m/s}} for vv in the equation p=mvp = mv to calculate the initial momentum of the clay 2, p2ix{p_{{\rm{2ix}}}}in the horizontal direction.

p2ix=(30g)(0m/s)=0kgm/s\begin{array}{c}\\{p_{{\rm{2ix}}}} = \left( {{\rm{30 g}}} \right)\left( {0{\rm{ m/s}}} \right)\\\\ = 0{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

Substitute p2iy{p_{{\rm{2iy}}}} for pp, 30g{\rm{30 g}} for mm, and 0m/s0{\rm{ m/s}} for vv in the equation p=mvp = mv to calculate the initial momentum of the clay 2, p2iy{p_{{\rm{2iy}}}}in the vertical direction.

p2iy=(30g)(2.0m/s)=(30g(103kg1g))(2.0m/s)=0.06kgm/s\begin{array}{c}\\{p_{{\rm{2iy}}}} = \left( {{\rm{30 g}}} \right)\left( {{\rm{2}}{\rm{.0 m/s}}} \right)\\\\ = \left( {{\rm{30 g}}\left( {\frac{{{{10}^{ - 3}}{\rm{ kg}}}}{{1{\rm{ g}}}}} \right)} \right)\left( {{\rm{2}}{\rm{.0 m/s}}} \right)\\\\ = 0.06{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

The total initial momentum is the sum of initial momentum of clay 1 and clay 2.

pi=p1i+p2i{p_{\rm{i}}} = {p_{{\rm{1i}}}} + {p_{{\rm{2i}}}}

Substitute 0.072kgm/s0.072{\rm{ kg}} \cdot {\rm{m/s}} for p1i{p_{{\rm{1i}}}} and 00 for p2i{p_{{\rm{2i}}}} in the equation pi=p1i+p2i{p_{\rm{i}}} = {p_{{\rm{1i}}}} + {p_{{\rm{2i}}}} to calculate the total initial momentum in horizontal direction pix{p_{{\rm{ix}}}}.

pix=0.072kgm/s+0=0.072kgm/s\begin{array}{c}\\{p_{{\rm{ix}}}} = 0.072{\rm{ kg}} \cdot {\rm{m/s}} + 0\\\\ = 0.072{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

Substitute 0kgm/s{\rm{0 kg}} \cdot {\rm{m/s}} for p1i{p_{{\rm{1i}}}} and 0.06kgm/s0.06{\rm{ kg}} \cdot {\rm{m/s}} for p2i{p_{{\rm{2i}}}} in the equation pi=p1i+p2i{p_{\rm{i}}} = {p_{{\rm{1i}}}} + {p_{{\rm{2i}}}} to calculate the total initial momentum in vertical direction, piy{p_{{\rm{iy}}}}.

piy=0+0.06kgm/s=0.06kgm/s\begin{array}{c}\\{p_{{\rm{iy}}}} = 0 + 0.06{\rm{ kg}} \cdot {\rm{m/s}}\\\\ = 0.06{\rm{ kg}} \cdot {\rm{m/s}}\\\end{array}

The total final momentum is the sum of final momentum of clay 1 and 2.

pfinal=p1f+p2f{p_{{\rm{final}}}} = {p_{1{\rm{f}}}} + {p_{2{\rm{f}}}}

Here, p1f{p_{1{\rm{f}}}} is the final momentum of the clay 1 and p2f{p_{{\rm{2f}}}} is the final momentum of clay 2.

Substitute m1vf{m_1}{v_{\rm{f}}} for p1f{p_{{\rm{1f}}}} and m2vf{m_2}{v_{\rm{f}}} for p2f{p_{2{\rm{f}}}} in the above equation pfinal=p1f+p2f{p_{{\rm{final}}}} = {p_{1{\rm{f}}}} + {p_{2{\rm{f}}}}.

pfinal=m1vf+m2vf{p_{{\rm{final}}}} = {m_1}{v_{\rm{f}}} + {m_2}{v_{\rm{f}}}

Substitute m1vf+m2vf{m_1}{v_{\rm{f}}} + {m_2}{v_{\rm{f}}} for pfinal{p_{{\rm{final}}}} in the conservation of momentum equation pinitial=pfinal{p_{{\rm{initial}}}} = {p_{{\rm{final}}}} and solve for vf{v_{\rm{f}}}.

pinitial=m1vf+m2vfvf=pinitialm1+m2\begin{array}{c}\\{p_{{\rm{initial}}}} = {m_1}{v_{\rm{f}}} + {m_2}{v_{\rm{f}}}\\\\{v_{\rm{f}}} = \frac{{{p_{{\rm{initial}}}}}}{{{m_1} + {m_2}}}\\\end{array}

Substitute vx{v_x} for vf{v_{\rm{f}}}, 0.072kgm/s0.072{\rm{ kg}} \cdot {\rm{m/s}} for pinitial{p_{{\rm{initial}}}}, 20g{\rm{20 g}} for m1{m_1}, and 30g{\rm{30 g}} for m2{m_2} in the equation vf=pinitialm1+m2{v_{\rm{f}}} = \frac{{{p_{{\rm{initial}}}}}}{{{m_1} + {m_2}}} and calculate vx{v_x} in the horizontal direction.

vx=0.072kgm/s20g+30g=0.072kgm/s50g(103kg1g)=1.44m/s\begin{array}{c}\\{v_x} = \frac{{0.072{\rm{ kg}} \cdot {\rm{m/s}}}}{{20{\rm{ g}} + 30{\rm{ g}}}}\\\\ = \frac{{0.072{\rm{ kg}} \cdot {\rm{m/s}}}}{{50{\rm{ g}}\left( {\frac{{{{10}^{ - 3}}{\rm{ kg}}}}{{1{\rm{ g}}}}} \right)}}\\\\ = 1.44{\rm{ m/s}}\\\end{array}

Substitute vy{v_y} for vf{v_{\rm{f}}}, 0.06kgm/s0.06{\rm{ kg}} \cdot {\rm{m/s}} for pinitial{p_{{\rm{initial}}}}, 20g{\rm{20 g}} for m1{m_1}, and 30g{\rm{30 g}} for m2{m_2} in the equation vf=pinitialm1+m2{v_{\rm{f}}} = \frac{{{p_{{\rm{initial}}}}}}{{{m_1} + {m_2}}} and calculate vy{v_y} in the vertical direction.

vy=0.06kgm/s20g+30g=0.06kgm/s50g(103kg1g)=1.2m/s\begin{array}{c}\\{v_y} = \frac{{0.06{\rm{ kg}} \cdot {\rm{m/s}}}}{{20{\rm{ g}} + 30{\rm{ g}}}}\\\\ = \frac{{0.06{\rm{ kg}} \cdot {\rm{m/s}}}}{{50{\rm{ g}}\left( {\frac{{{{10}^{ - 3}}{\rm{ kg}}}}{{1{\rm{ g}}}}} \right)}}\\\\ = 1.2{\rm{ m/s}}\\\end{array}

Use the magnitude equation to calculate speed.

Substitute 1.44m/s1.44{\rm{ m/s}} for vx{v_x} and 1.2m/s1.2{\rm{ m/s}} for vy{v_y} in the equation v=vx2+vy2v = \sqrt {v_x^2 + v_y^2} .

v=(1.44m/s)2+(1.2m/s)2=1.87m/s\begin{array}{c}\\v = \sqrt {{{\left( {1.44{\rm{ m/s}}} \right)}^2} + {{\left( {1.2{\rm{ m/s}}} \right)}^2}} \\\\ = 1.87{\rm{ m/s}}\\\end{array}

Substitute 1.44m/s1.44{\rm{ m/s}} for vx{v_x} and 1.2m/s1.2{\rm{ m/s}} for vy{v_y} in the equation θ=arctan(vyvx)\theta = \arctan \left( {\frac{{{v_y}}}{{{v_x}}}} \right).

θ=arctan(1.2m/s1.44m/s)=39.8\begin{array}{c}\\\theta = \arctan \left( {\frac{{1.2{\rm{ m/s}}}}{{1.44{\rm{ m/s}}}}} \right)\\\\ = 39.8^\circ \\\end{array}

Ans:

The magnitude of the final speed of the resulting 50 g clay is 1.87 m/s.

The direction of the final speed of the resulting 50 g clay is 39.839.8^\circ north of east.